Two cross country runners are 440 yards apart and are running toward each other, one at 8mph and the other at 10 mph. In how many seconds will they meet?
Answer by checkley71(8403)

440 YDS=.25 MILES
8+10=18 MPH
DISTANCE=RATE*TIME
.25=18*T
T=.25/18
T=.0138888 HOURS OR .013888*3600=50 SECONDS ANSWER..